Example: Lets say that the managers for the manufacturing plant where the Ford Edge is made have predicted $36 million of MOH and 960,000 DL hours for the year. They have also estimated direct labor for the year to cost $24 million *Compute the predetermined MOH rate for the year if DL hours is used as the allocation base. Now lets say Job #204(batch of 15 identical vehicles) took 740 DL hours to complete and that the assembly line workers who worked on the job were paid $22 per hour. Furthermore, S150,000 of DM was requisitioned for the job How much MOH should be allocated to Job #204? . What is the total cost of Job #2042
Answer #1

Predetermined MOH rate for the year = $360,00,000/9,60,000

MOH rate per labor hour = $37.5

MOH to be allocated to Job#204 = 37.5*740 = $27,750

Computation of total cost of job#204:

Direct material = $1,50,000

Direct labor cost = 22*740 hours = $16,280

MOH allocated to the job = $27,750

Total cost of job#204 = $1,94,030.

Thanks,

Let me know if you have any doubts.